### v1

# Residual lifetime prediction for heterogeneous degradation data by Bayesian semi-parametric method

Abstract
Degradation data are considered for assessing reliability in highly reliable systems. The usual assumption is that degradation units come from a homogeneous population. But in presence of high variability in the manufacturing process, this assumption is not true in general; that is different sub-populations are involved in the study. Predicting residual lifetime of a functioning unit is a major challenge in the degradation modeling especially in heterogeneous environment. To account for heterogeneous degradation data, we have proposed a Bayesian semi-parametric approach to relax the conventional modeling assumptions. We model the degradation path using Dirichlet process mixture of normal distributions. Based on the samples obtained from posterior distribution of model parameters we obtain residual lifetime distribution for individual unit. Transformation based MCMC technique is used for simulating values from the derived residual lifetime distribution for prediction of residual lifetime. A simulation study is undertaken to check performance of the proposed semi-parametric model compared with parametric model. Fatigue Crack Size data is analyzed to illustrate the proposed methodology.
keywords:
Degradation, General path model, Residual Life time, Bayesian semi-parametric, Dirichlet Process.
1 Introduction
For some highly reliable engineering products, it is difficult to estimate reliability due to the fact that products take too long time to fail. In this context, degradation measures taken over time are used to estimate reliability. The evolution of the degradation measures may be observed using sensor technology through a procedure known as Condition Monitoring (See Nelson [1]). Some examples of degradation data include vibration signals for monitoring excessive wear induced in rotating machinery, acoustic emissions for monitoring crack propagation, temperature changes and oil debris for engine lubrication, decrease of brightness of light bulbs, etc. Inferences on lifetime distribution and residual lifetime of a product can be done by modeling the underlying degradation mechanism that represents the evolution of degradation resulting failure. There have been a number of works on degradation modeling, see for example Lu and Meeker [2], Padgett and Tomlinson [3], Muller and Zhang [4] and Park and Padgett [5]. Degradation data provide more information than the bare lifetimes and hence precise inferences can be made as discussed in Lu and Meeker [2]. The modeling of degradation data can be considered by different processes. The commonly used degradation processes are the Weiner, Inverse-Gaussian, and Gamma processes (See Lawless and Crowder [6]) including general path model. The general path model, introduced by Lu and Meeker [2] is widely used to model degradation data. Ye and Xie [7] have done a comparative study between general path model and stochastic processes for modeling degradation data.
Estimation of residual lifetime distribution of systems operating in the field plays a key role in implementing condition-based maintenance decision making (see Jardine and Banjevic [8]). Different degradation models are considered to derive the distribution function of residual lifetime. Gebraeel et al. have introduced ([9], [10]) a Bayesian degradation path model for predicting residual life time of units. Zhou et al. [11] have proposed an empirical Bayes approach to update the stochastic parameters of the degradation model for predicting soft-failure of a functioning device. Liu et al. [12] have considered modified Weiner process to predict the remaining useful lifetime; and Si et al. [13] have introduced a Wiener-process-based degradation model with a recursive filter algorithm for remaining useful lifetime prediction. A key assumption in most of the degradation models is that the degradation rate is homogeneous, see for example, Lu and Meeker [2] and Robinson and Crowder [14]. However, in practice, the assumption of homogeneity is not appropriate, particularly in presence of high variability in the manufacturing process, where the population consists of a number of homogeneous sub-populations. Ye et al. [15] have considered heterogeneous degradation rates for estimating the distribution of remaining useful life. Wen et al. [16] considered Wiener process to predict remaining lifetime in presence of heterogeneous population.
Different mixture models are used for modeling heterogeneous failure time data. Mixture of gamma distributions and mixture of normal distributions (Kontar et al. [17]) are commonly used for mixture models. In the context of degradation data, Yuan and Ji [18] analyzed Laser emitters data and noticed that some units degrade faster than other units indicating the fact that they come from different sub-populations. They have considered a finite mixture of normal distributions for random effects in the linear degradation model. Model selection criteria like Akaike information criterion (AIC) and Bayesian information criterion (BIC) are used to estimate the number of mixture components. One of the major difficulty in using finite mixture models is defining the true number of components in the mixture, which denotes different sub-population under study. Limiting true number of mixture components may lead to wrong estimate of parameters. To overcome the above limitation of pre-specifying the number of sub-populations, Bayesian nonparametric models are considered in recent years.
Dirichlet process introduced by Fergusan (See Fergusan [19]) is one of the most popular Bayesian nonparametric methods. Dirichlet process prior puts probability 1 to the set of discrete probability measures. To get rid of this problem Dirichlet process mixture (DPM) of normal distributions is used in the literature (See Escobar and West [20]), when the true distribution is continuous. Lo [21] introduced Monte Carlo simulation based methods for density estimation using Dirichlet process mixtures. Escobar ([22], [20]) and MacEachern [23] developed Gibbs sampling methods for Dirichlet process in a normal mixture model. But these techniques tended to produce limited posterior inference for example the Markov chain produced by these techniques tends to mix very slowly. To overcome this problem, Ishwaran and Zarepour ([25], [24]) introduced Gibbs sampling methods for the approximate Dirichlet process which consider a truncation approximation of Dirichlet process. In the context of modeling heterogeneous degradation data, Bayesian non-parametric method have gained popularity in recent times, see for example see Santos and Loschi [26], Li et al. [27] and Cheng and Yuan [28]. Nguyen et al. [29] used Bayesian non-parametric model for remaining useful life prediction of individual units with sparse degradation data.
In this article, we introduce a simulation based prediction method for residual lifetime of a functioning unit, when the unit comes from a heterogeneous population. We estimate residual lifetime distribution of a functioning unit conditioned on the historical degradation measures of other units and degradation observation of the concerned unit. Firstly, we consider a Bayesian semi-parametric technique, where the degradation path is modeled using general path model and the distribution of random effect is modeled by Dirichlet process mixture of normal distribution, and parametric prior is considered for other model parameters. Utilizing the idea of Ishwaran et al. [24], we propose a hierarchical model and generate samples from posterior distributions of parameters by Gibbs sampling method. We show that the residual lifetime distribution depends on the posterior samples of parameters and it does not have closed form expression. So an approximate residual lifetime distribution is derived and finally transformation based MCMC technique [30] is used for simulating samples from the approximated distribution. Finally, we use the simulated samples for predicting residual life time of a unit. We compare the accuracy of predicted results produced by our proposed Bayesian semi-parametric method with Bayesian parametric method, where a parametric distribution is considered for the random effect and other parameters.
The rest of this paper is organized as follows. Degradation model is discussed in Section 2. Residual lifetime distribution is obtained based on proposed degradation model in Section 3. We discuss method of generating observations from posterior distribution in Section 4. A simulation study is undertaken to asses the performance and efficiency of the proposed Bayesian semi-parametric method in Section 5. A real-life data on fatigue crack size is analyzed to illustrate the proposed methodology in Section 6. We conclude this article in section 7.
2 Degradation model
Consider the situation where degradation measurements are taken for different units at some fixed time points and the units are from a heterogeneous population where degradation rate vary significantly among the units. Appropriate modeling of degradation data is required in this heterogeneous situation.
2.1 Bayesian semi-parametric degradation model
Let $Y_{t}$ be random variable for degradation measurement of a unit at the $t$th time point. Consider the general path model introduced by Meeker [2], where $Y_{t}$ is assumed to have the following representation.
$Y_{t}=\eta(t;\bm{\alpha},\bm{\beta})+\epsilon_{t}\hskip 11.38109pt\forall\hskip 2.27621ptt\geq 0,$ (1)
where $\eta(\cdot;\bm{\alpha},\bm{\beta})$ denotes the true degradation path, $\bm{\alpha}=(\alpha_{1},\alpha_{2},.....\alpha_{k_{1}})^{t}$ is a $k_{1}$ 1 vector of parameters that are common to all units, known as fixed effect and $\bm{\beta}=(\beta_{1},\beta_{2},.....,\beta_{{k_{2}}})^{t}$ is a $k_{2}$ 1 vector of random effects for representing individual unit characteristics. Further it is assumed that the measurement errors $\epsilon_{t}$ are i.i.d $\mathcal{N}(0,\sigma_{\epsilon}^{2})$, $\forall\hskip 2.27621ptt$. It is assumed that the random effect $\bm{\beta}$ is independent of $\epsilon_{t}$, for $\forall\hskip 2.27621ptt\geq 0$. Suppose we have $n$ sample units randomly selected from the population and degradation measurements are taken at time points $t_{i1},t_{i2},...,t_{in_{i}}$ for $i$th individual where $i=1,......n$. Let $Y_{ij}$ be random variable for the degradation measurement of $i$th unit at time $t_{ij}$. Using equation $(1)$, $Y_{ij}$ is represented as follows.
[…]
For each sample unit $i$, the degradation path is defined to be an observed sequence of degradation readings $\textbf{y}_{i}$ over time $\textbf{t}_{i}$, where $\textbf{y}_{i}=(y_{i1},...,y_{i{n_{i}}})$, $\textbf{t}_{i}=(t_{i1},...,t_{in_{i}})$ and $\bm{\beta}_{i}$ is the random effect for unit $i$. Linear and nonlinear structures for true degradation path in general path models are considered in different contexts. For example Lu and Meeker [2] have considered a non-linear model to estimate reliability of a unit. They considered that a unit fails when the true path $\eta(\cdot)$ crosses some predetermined threshold value, say $D$. On the other hand, Robinson and Crowder [14], Zhou et al. [11], Gebraeel et al. [10], assumed that failure of a unit occurs if its observed degradation measurement reaches a predetermined threshold $D$. This threshold value may be fixed or may vary with time.
One of the important issues in using the general path models is the specification of the distribution of the random effects. The random-effects $\bm{\beta}$ represents unit wise effect and can be interpreted as degradation rates. Most of the existing literature assumes that the units under test originate from homogeneous populations, hence a unimodal distribution is assumed for the random effects for modeling degradation path. But in practice, it may happen that units come from heterogeneous population consisting of different homogeneous sub-populations. In this context, Yuan and Ji [18] have used finite mixture distribution for random effects to incorporate heterogeneity. Bayesian method is considered to account for uncertainty of number of mixture components. But in finite mixture modeling, one of the major problems is that the model assumes a finite mixture of $K$ component regardless of sample size $n$ and that ignores the fact that $K$ can also grow as the sample size $n$ increases. It may happen that a unit is degrading significantly different from all other units and a finite mixture model will not able to recognize this unit as it is coming from distinct sub-population other than $K$ many components chosen for modeling. This may lead to biased parameter estimates which in turn effects the estimation of reliability and residual lifetime distribution. To deal with this situation infinite mixture model can be considered (for example see Santos and Loschi [26]), where the number of mixture components slowly grows with sample sizes. A Bayesian alternative is to consider unknown distribution function $F$ for the corresponding random variable and hence construction of a prior for $F$ is necessary. This will also help to avoid misspecification of the distribution. In this article we consider a unknown distribution function for random effects to overcome the above-mentioned problems.
Let us consider $\textbf{Y}_{train}$ = $\{Y_{ij}\}_{i=1,...,n}^{j=1,...,n_{i}}$ as the random variable that denotes the degradation measures of $n$ units. We consider these $n$ units as training set. Suppose that there is a new unit with degradation observations at some time points $t_{1},....,t_{k}$, where $t_{1}<....<t_{k}$. Let us consider the new unit as $(n+1)$th unit, where $\textbf{Y}_{n+1}=\textbf{Y}_{new}=\{Y_{new,t_{1}},....,Y_{new,t_{k}}\}$ is the random variable for degradation measurements of the new unit. Let us consider $\{y_{new,t_{1}},....,y_{new,t_{k}}\}$ as the observed degradation measures of the new unit and $\textbf{t}_{n+1}=\{t_{1},...,t_{k}\}$. We consider the random variable $\bm{Y}=[\textbf{Y}_{train},\textbf{Y}_{new}]$, which represents the degradation measures of the units in the training set and the new unit. We utilize the entire degradation information contained in the $(n+1)$ units to model the degradation path which helps in deriving the residual lifetime distribution of the new unit. In this work we use mixed effect linear model introduced by Lu and Meeker [2] for degradation modeling with one fixed effect $\alpha$ and random effect $\beta_{i}$, $i=1,...(n+1)$. Let us consider $\beta_{(n+1)}$ be the random effect on this new unit, which we represent as $\beta_{new}$. We assume that the random effect $\beta_{i}|F\overset{\mathrm{iid}}{\sim}F$, $i=1,...,(n+1)$, where $F$ is unspecified distribution and has a density function. The proposed mixed effect model is given as:
$Y_{ij}$ $=\alpha+\beta_{i}t_{ij}+\epsilon_{ij},\hskip 5.69054pti=1,....,(n+1),j=1,....,n_{i}$ (3)
[…]
Next we consider selection of prior for $F$. One of the commonly used priors for $F$ is the Dirichlet process (DP) prior that is a random probability measure defined on the space of distribution functions, introduced by Fergusan [19]. Dirichlet process prior is almost surely (a.s) discrete hence it would be inappropriate to use this since we assumed that $F$ has a density function. In order to overcome this problem, we consider Dirichlet process mixture (DPM) of continuous distributions (See Escobar [20], Lo [21]). Also we consider a prior $\pi_{1}$ on the fixed effect parameter $\alpha$ and the $\pi_{2}$ on variance of measurement error $\sigma_{\epsilon}^{2}$ and they are independent to each other. In the proposed method, the degradation path is modelled using general path model and distribution of random effect is modelled through Bayesian non-parametric method, while parametric prior is considered for other model parameters. So essentially we consider a Bayesian semi-parametric technique for our analysis.
In a Dirichlet process mixture modeling, the distribution function of random effect is represented by mixture over some simple parametric distribution functions where the mixing distribution is given by a Dirichlet process prior. We consider that $\beta_{i}|\theta_{i}\sim f(\cdot)$, $i=1,...,(n+1)$, where $f(\cdot)$ is a continuous distribution on $\mathbb{R}$ and $\theta_{i}$’s are unobserved random elements. Suppose that $\theta_{i}|G\overset{\mathrm{iid}}{\sim}G$ and G has a Dirichlet process prior denoted by $G\sim DP(\gamma,G_{0})$, where $\theta_{i}$in $\Theta$, $i=1,...,(n+1)$ and $G_{0}$ is the center or baseline probability measure on the measurable space ($\Theta$$\mathscr{B}$), where $\mathscr{B}$ is the corresponding Borel $\sigma$-algebra and $\gamma\in R^{+}$, considered as concentration parameter. Under these assumptions, for $i=1,.....,(n+1)$, and $j=1,....,n_{i}$, the proposed degradation model can be hierarchically represented as :
$Y_{ij}|\alpha,\beta_{i},\sigma_{\epsilon}^{2}\overset{\mathrm{ind}}{\sim}\mathcal{N}(\alpha+\beta_{i}t_{ij},\sigma_{\epsilon}^{2})$ (4)
${\beta}_{i}|\theta_{i}\overset{\mathrm{ind}}{\sim}f(\cdot)$
$\theta_{i}|G\overset{\mathrm{iid}}{\sim}G$
$G\sim DP(\gamma,G_{0})$
$\alpha\sim\pi_{1}(\alpha)$
$\sigma_{\epsilon}^{2}\sim\pi_{2}(\sigma_{\epsilon}^{2})$
Note that the parameter $\gamma$ controls the concentration of the prior for G about $G_{0}$. It can be proved that, for any measurable subset $A$ of $\Theta$, $E[G(A)]={G_{0}}(A)$ and $Var[G(A)]=\frac{{G_{0}}(A)[1-{G_{0}}(A)]}{1+\gamma}$. Observe that $\forall A\in\mathscr{B}$, $G(A)$ is highly concentrated about $G_{0}(A)$ for large values of $\gamma$. On the other hand, as $\gamma$ tends to zero the expected shape of G is different from the one assumed by the baseline probability measure $G_{0}$. Since $G$ is a discrete probability measure with probability 1, so there will be repetition among the drawn values from this distribution, in other words if we draw $m$ values say $\theta_{1},.....,\theta_{m}$ from $G$, then there will be $k_{m}$ unique values of $\theta_{i}$’s where $k_{m}\leq m$. One can observe that clustering is induced among $\beta_{i}$’s, $i=1,...,(n+1)$ due to the fact that there are similar values among the unobserved random element $\theta_{i}$’s. So there is a positive probability that given $\theta_{i}$’s the distribution of random effect $\beta_{i}$’s are same for distinct $i$’s. The expected number of clusters or, equivalently, the number of distinct $\theta_{i}$ among the $m$ drawn sample values is given by $E[k_{m}]=\sum_{i=1}^{m}\frac{\gamma}{\gamma+i-1}$. If $m\rightarrow\infty$, it follows that $E[k_{m}]\approx\gamma\log(\frac{\gamma+m}{\gamma})$. Our next goal is to choose a suitable expression for $f(\cdot),\pi_{1}(\cdot)$ and $\pi_{2}(\cdot)$ and construct the hierarchy of the model.
2.2 Dirichlet process mixture of normal distribution for degradation model
As discussed in Section 2.2, we assume that the mixing probability measure $G$ has a Dirichlet process prior denoted by $G\sim DP(\gamma,G_{0})$. Sethuraman [32] showed that $G$ can be represented as, $G(\cdot)=\sum\limits_{h=1}^{\infty}p_{h}\delta_{m_{h}}(\cdot)$, where $m_{h}$ are random samples from $G_{0}$, $0\leq p_{h}\leq 1$ $\forall$ $h$ and $p_{1}=V_{1}$, $p_{h}=(1-V_{1})(1-V_{2})...(1-V_{h-1})V_{h}$ $\forall$ $h\geq 2$. Ishawarn and Zarepour [25] represented this sum of infinitely many terms by sum of finitely many $N$ terms. This method approximate Dirichlet process prior by $\mathbb{G}_{N}$, where $\mathbb{G}_{N}(\cdot)=\sum\limits_{h=1}^{N}p_{h}\delta_{Z_{h}}(\cdot)$ and $\delta_{Z}(\cdot)$ denotes a discrete measure concentrated at $Z$. Note that $\mathbb{G}_{N}$ is a random probability measure, and as $N\rightarrow\infty$, it converges almost surely to a Dirichlet process with baseline probability measure $G_{0}$ and Dirichlet mass parameter $\gamma$, denoted by $DP(\gamma,G_{0})$. Let us consider, $\textbf{Z}=(Z_{1},..,Z_{N})$, and $Z_{1},....,Z_{N}$ are i.i.d random variables with distribution $G_{0}$ and also they are independent of $\textbf{p}=(p_{1},...,p_{N})$. A probability measure simulated from the prior $\mathbb{G}_{N}$ is defined by choosing its random weights $p_{1},...,p_{N}$ by the stick-breaking construction, where $p_{1}=V_{1}$, $p_{h}=(1-V_{1})(1-V_{2})...(1-V_{h-1})V_{h},\hskip 2.27621pth=2,...,N$, with $V_{1},V_{2},...V_{N-1}$ are i.i.d $Beta(1,\gamma)$ random variables, setting $V_{N}=1$ ensures that $\sum\limits_{h=1}^{N}p_{h}=1$.
Ishawaran and Zarepour [25] utilized approximated Dirichlet process $\mathbb{G}_{N}$ and introduced block Gibbs sampler to do direct inference for $\mathbb{G}_{N}$. They have constructed an efficient MCMC method, which recast the nonparametric hierarchical model completely in terms of random variables. Let $K_{1},......,K_{n+1}$ be the classification variables of the random effects $\beta_{1},......,\beta_{n+1}$ and $K_{i}$’s are conditionally independent random variables given p which identifies $Z_{k}$ with each associated $\theta_{i}$, particularly $\theta_{i}=Z_{K_{i}}$, $i=1,...,(n+1)$. The clustering nature of the hidden variables $\theta_{i}$’s are described by the classification variables $K_{1},......,K_{n+1}$. Let us consider $\pi_{3}(\cdot)$ and $\pi_{4}(\cdot)$ be the prior distributions for p and Z respectively. It follows that the hierarchy for distribution of $\beta_{i}$. $i=1,....,(n+1)$ as presented in equation (4) can be rewritten as (See Ishawaran and Zarepour [25])
$\beta_{i}|\textbf{Z},K_{i}\sim f(\beta_{i}|Z_{K_{i}})$ (5)
$K_{i}|\textbf{p}\sim\sum\limits_{h=1}^{N}p_{h}\delta_{h}(\cdot)$
$\sim\pi_{3}(\textbf{p})$ p )
$\sim\pi_{4}(\textbf{Z})$ Z )
This representation also helps to the fact that any degradation trajectory that is significantly different from the others can be classified as a distinct cluster. In this work we have taken the truncation parameter $N$ as the total number of units on which the proposed degradation model is built.
We consider a normal distribution as a prior for the fixed effect parameter $\alpha$ and hierarchical priors are considered for its parameters. To complete the prior specification of $\mathbb{G}_{N}(\cdot)$, we consider $Z_{h}=(\mu_{h},\sigma_{h}^{2})$, where $\mu_{h}$ and $\sigma_{h}^{2}$ are mean and variance of conditional distribution of random effect given $\bm{\mu},\bm{\sigma}^{2},K_{i}=h$ which is normal and denoted as $\mathcal{N}(\cdot)$ in this work. Normal prior is considered for $\mu_{h}$’s and gamma prior for $(\sigma_{h}^{2})^{-1}$, $h=1,....,N$ which we denote as $\mathcal{G}(\cdot)$. We consider Gamma distribution with the form $\mathcal{G}(a_{0},b_{0})$ so that mean is $\frac{a_{0}}{b_{0}}$. Gamma prior is considered for inverse of variance of error measurements $\sigma_{\epsilon}^{2}$. We also assume that $\mu_{h}\gg\sigma_{h}$, $h=1,...,N$. We discuss in details on the prior specification in Section 4. Observe that, classification variables of random effects $K_{i}$’s, $i=1,...,(n+1)$ are conditionally independent random variables given p, where p is constructed by stick-breaking construction. Note that $K_{n+1}$ is the classification variable corresponding to random effect of the new unit. The hierarchy of the model is given below :
[…]
$\hskip 113.81102pt\alpha\sim\mathcal{N}(\mu_{\alpha},\sigma_{\alpha}^{2})$
$\hskip 113.81102pt\mu_{\alpha}\sim\mathcal{N}(\mu_{1},\sigma_{1}^{2})$
$\hskip 108.12047pt(\sigma_{\alpha}^{2})^{-1}\sim\mathcal{G}(a_{1},a_{2})$
$\hskip 99.58464pt\beta_{i}|\bm{\mu},\bm{\sigma}^{2},K_{i}\sim\mathcal{N}(\mu_{K_{i}},{\sigma}_{K_{i}}^{2}),\hskip 5.69054pti=1,......,n+1$
$\hskip 99.58464ptK_{i}|p\sim\sum\limits_{h=1}^{N}p_{h}\delta_{h}(\cdot),\hskip 5.69054pti=1,......,n+1$
$\hskip 71.13188ptp_{1}=V_{1},\hskip 5.69054ptp_{h}=V_{h}\prod\limits_{j=1}^{h-1}(1-V_{j}),\hskip 5.69054pth=2,...,N,\hskip 5.69054ptV_{l}|\gamma\overset{\mathrm{i.i.d}}{\sim}Beta(1,\gamma),$
$\hskip 71.13188pt\hskip 5.69054ptl=1,....,N-1\hskip 5.69054pt\&\hskip 5.69054ptV_{N}=1$
$\hskip 113.81102pt\mu_{h}|\sigma_{z}^{2}\sim\mathcal{N}(m_{\mu},\sigma_{z}^{2}),\hskip 5.69054pth=1,......,N$
$\hskip 113.81102pt({\sigma_{z}^{2}})^{-1}\sim\mathcal{G}(\tau_{1},\tau_{2})$
$\hskip 113.81102pt({\sigma_{h}^{2}})^{-1}\sim\mathcal{G}(a,b),\hskip 5.69054pth=1,......,N$
$\hskip 113.81102pt(\sigma_{\epsilon}^{2})^{-1}\sim\mathcal{G}(a_{\epsilon},b_{\epsilon})$
$\hskip 119.50157pt\gamma\sim\mathcal{G}(\eta_{1},\eta_{2})$
The posterior distribution of the parameters is given by
$\pi(\bm{\beta},\alpha,\mu_{\alpha},\sigma_{\alpha}^{2},\bm{\mu},\sigma_{z}^{2},\textbf{K},\textbf{p},\bm{\sigma}^{2},\sigma_{\epsilon}^{2},\gamma|\bm{Y}=\bm{y}),$ (7)
where $\bm{\mu}=(\mu_{1},.......,\mu_{N})$, $\bm{\sigma}^{2}=(\sigma_{1}^{2},.......,\sigma_{N}^{2})$, $\bm{\beta}=(\beta_{1},.....,\beta_{n+1})$, $\textbf{K}=(K_{1},......,K_{n+1})$ and $\textbf{p}=(p_{1},......,p_{N})$. In the next section we consider derivation of residual lifetime distribution of a unit conditioned on the observed degradation measurements. We show that this distribution depends on the samples from posterior distribution of parameters and hence it is required to generate samples from the above-mentioned posterior distribution presented in (7).
3 Residual lifetime distribution
Our goal is to derive distribution function of residual lifetime of a new unit for which we have degradation measurements at time points $t_{1},....,t_{k}$. Suppose the degradation observations of the new unit has not crossed threshold value $D$ till time $t_{k}$, that is $\underset{1\leq a\leq k}{max}\{y_{new,t_{a}}\}$ $\leq D$. The residual lifetime $T_{new}$ of new unit functioning at time $t_{k}$ is defined to be the first time when the degradation measurement reaches the threshold value $D$ subtracted from $t_{k}$, which is expressed as follows.
$T_{new}=$ $\underset{s_{1}\in(t_{k},\infty)}{\text{inf}}\{s_{1}-t_{k}:Y_{new,s_{1}}\geq D\}$ (8)
$=$ $\underset{s\in(0,\infty)}{\text{inf}}\{s:Y_{new,s+t_{k}}\geq D\},\hskip 14.22636pts=s_{1}-t_{k},$
Next we derive the distribution of $T_{new}$. We make the following assumptions for deriving the residual lifetime distribution of a functioning unit.
  1. 1.
    The random variable for residual lifetime has support on $(0,\infty)$.
  2. 2.
    Once the degradation path exceeds the predefined threshold $D$ for first time say at time point $t^{*}$ it will never crosses back the threshold after $t^{*}$, that is there does not exist $y\geq 0$ such that $P(Y_{t^{*}+y}\leq D)>0$.
The distribution function of $T_{new}$ is obtained as.
$F_{T_{new}}(t)=$ $P[T_{new}\leq t]$ (9)
$=$ $1-P[T_{new}>t]$
$=$ $1-P[\underset{s\in(0,\infty)}{\text{inf}}(s:Y_{new,s+t_{k}}\geq D)>t]$
$=$ $2\displaystyle 1-P[Y_{new,t+t_{k}}<D],\hskip 5.69054ptby\hskip 5.69054pt\text{assumption}\hskip 2.27621pt$ 2
$=$ $P[Y_{new,t+t_{k}}\geq D].$
We compute distribution function of random variable $T_{new}$ given $\bm{Y}=\bm{y}$, which is denoted as $F_{T_{new}|\bm{Y}=\bm{y}}(t)$ where $\bm{y}$ is the observed value of the random variable $\bm{Y}$. This implies that it is enough to derive $P[Y_{new,t+t_{k}}\geq D|\bm{Y}=\bm{y}]$, since the event $[T_{new}\leq t]$ equals the event $[Y_{new,t+t_{k}}\geq D]$ under the assumption 2. Thus it is also possible to derive residual lifetime distribution of each unit of the training set by considering the individual unit as the unit for which residual lifetime is to be predicted and rest of the units as training set. So, if there are $n$ units in training set, then we will fix one unit as new unit and use $(n-1)$ units as training data. This process can be done for all $n$ units. In this work we consider that failure of a unit occurs when its observed degradation measurement crosses a predetermined threshold value say $D$. Sometimes degradation measurements of a unit are available after its degradation path crosses the threshold value. In our proposed method these information are also used to compute residual lifetime distribution. The residual lifetime distribution of a new unit conditioned on $\bm{Y}=\bm{y}$, is given by:
$F_{T_{new}|\bm{Y}=\bm{y}}(t)=$ $P[T_{new}\leq t|\bm{Y}=\bm{y}]$
$=$ $P[Y_{new,t+t_{k}}\geq D|\bm{Y}=\bm{y}]$
$=$ $\int_{\bm{\gamma_{1}}}P[Y_{new,t+t_{k}}\geq D,\bm{\gamma_{1}}|\bm{Y}=\bm{y}]d\bm{\gamma_{1}},\hskip 2.84526pt\bm{\gamma_{1}}=(\alpha,\beta_{1},....,\beta_{n},\beta_{new},\sigma_{\epsilon}^{2})$
[…]
$=$ $E_{\bm{\gamma_{1}}|\bm{Y}=\bm{y}}\big{[}P[Y_{new,t+t_{k}}>D|\bm{\gamma_{1}}]\big{]}.\hskip 176.40707pt(10)$
It may be noted that $F_{T_{new}|\bm{Y}=\bm{y}}(t)$ can not be obtained analytically. We compute it approximately based on observations generated from the posterior distribution of $\bm{\gamma_{1}}$. Let $\alpha^{(i)}$, $\beta_{1}^{(i)}$,…., $\beta_{n}^{(i)}$, $\beta_{new}^{(i)}$ and $(\sigma_{\epsilon}^{2})^{(i)}$, $i=1,.....,N_{1}$ are the samples generated from posterior distribution of $\bm{\gamma_{1}}$ and $N_{1}$ is number of posterior samples. We denote this approximated distribution function as $F_{T_{new}|\bm{Y}=\bm{y}}^{approx}(t)$, which is given as follows.
$F_{T_{new}|\bm{Y}=\bm{y}}^{approx}(t)$ $=$ $\frac{1}{N_{1}}\sum_{i=1}^{N_{1}}P[Y_{new,t+t_{k}}\geq D|\bm{\gamma_{1}}=\bm{\gamma_{1}}^{(i)}]$
$=$ $\frac{1}{N_{1}}\sum_{i=1}^{N_{1}}P[Y_{new,t+t_{k}}\geq D|\alpha=\alpha^{(i)},\beta_{1}=\beta_{1}^{(i)},...,\beta_{n}=\beta_{n}^{(i)},\beta_{new}=\beta_{new}^{(i)},\sigma_{\epsilon}^{2}=(\sigma_{\epsilon}^{2})^{(i)}]$
$=$ $\frac{1}{N_{1}}\sum_{i=1}^{N_{1}}P[Y_{new,t+t_{k}}\geq D|\alpha=\alpha^{(i)},\beta_{new}=\beta_{new}^{(i)},\sigma_{\epsilon}^{2}=(\sigma_{\epsilon}^{2})^{(i)}]$
$=$ $\frac{1}{N_{1}}\sum_{i=1}^{N_{1}}P[\alpha^{(i)}+\beta_{new}^{(i)}(t+t_{k})+\epsilon_{new,t+t_{k}}\geq D|\alpha=\alpha^{(i)},\beta_{new}=\beta_{new}^{(i)},\sigma_{\epsilon}^{2}=(\sigma_{\epsilon}^{2})^{(i)}]$
$=$ $\frac{1}{N_{1}}\sum_{i=1}^{N_{1}}P[\epsilon_{new,t+t_{k}}\geq D-\alpha^{(i)}-\beta_{new}^{(i)}(t+t_{k})|\alpha=\alpha^{(i)},\beta_{new}=\beta_{new}^{(i)},\sigma_{\epsilon}^{2}=(\sigma_{\epsilon}^{2})^{(i)}]$
$=$ $\frac{1}{N_{1}}\sum_{i=1}^{N_{1}}P[\epsilon_{{new,t+t_{k}}}\geq D-\alpha^{(i)}-\beta_{new}^{(i)}(t+t_{k})|\sigma_{\epsilon}^{2}=(\sigma_{\epsilon}^{2})^{(i)}]\hskip 5.69054pt$
$=$ $\frac{1}{N_{1}}\sum_{i=1}^{N_{1}}P\Big{[}\frac{\epsilon_{new,t+t_{k}}}{\sigma_{\epsilon}^{(i)}}\geq\frac{D-\alpha^{(i)}-\beta_{new}^{(i)}(t+t_{k})}{\sigma_{\epsilon}^{(i)}}|\sigma_{\epsilon}^{2}=(\sigma_{\epsilon}^{2})^{(i)}\Big{]}$
[…]
where $\Phi(\cdot)$ is cdf of standard normal random variable. So we approximate the residual lifetime distribution by the expression derived in equation (11). For the approximated distribution we have,
$F_{T_{new}|\bm{Y}=\bm{y}}^{approx}(0)=$ $P[T_{new}\leq 0|\bm{Y}=\bm{y}]$ (12)
$=$ $\frac{1}{N_{1}}\sum_{i=1}^{N_{1}}\Phi\Big{[}\frac{\alpha^{(i)}+\beta_{new}^{(i)}t_{k}-D}{\sigma_{\epsilon}^{(i)}}\Big{]}>0.$
So, equation (12) implies that the random variable representing the residual lifetime of a new unit given the entire degradation data can take negative values. But we assumed that the support of $T_{new}$ is $(0,\infty)$. Because of this, we consider the distribution function for $T_{new}$ given $\bm{Y}=\bm{y}$ with an additional constraint $T_{new}>0$ which we denote as $F_{T_{new}|\bm{Y}=\bm{y},T_{new}>0}(t)$ and $F_{T_{new}|\bm{Y}=\bm{y},T_{new}>0}^{approx}(t)$ is the corresponding approximated distribution function. This expression of this distribution function is given as below:
$F_{T_{new}|\bm{Y}=\bm{y},T_{new}>0}(t)=$ $P[T_{new}\leq t|T_{new}>0,\bm{Y}=\bm{y}]$ (13)
$=$ $\frac{P{(T_{new}\leq t,T_{new}>0|\bm{Y}=\bm{y})}}{P{(T_{new}>0|\bm{Y}=\bm{y})}}$
$=$ $\frac{P{(0<T_{new}\leq t|\bm{Y}=\bm{y})}}{P{(T_{new}>0|\bm{Y}=\bm{y})}}$
$=$ $\frac{P{(T_{new}\leq t|\bm{Y}=\bm{y})-P(T_{new}<0|\bm{Y}=\bm{y})}}{1-P{(T_{new}\leq 0|\bm{Y}=\bm{y})}}$
$F_{T_{new}|\bm{Y}=\bm{y},T_{new}>0}^{approx}(t)=$ $\frac{\frac{1}{N_{1}}\sum_{i=1}^{N_{1}}\Phi\Big{[}\frac{\alpha^{(i)}+\beta_{new}^{(i)}(t+t_{k})-D}{\sigma_{\epsilon}^{(i)}}\Big{]}-\frac{1}{N_{1}}\sum_{i=1}^{N_{1}}\Phi\Big{[}\frac{\alpha^{(i)}+\beta_{new}^{(i)}t_{k}-D}{\sigma_{\epsilon}^{(i)}}\Big{]}}{1-\frac{1}{N_{1}}\sum_{i=1}^{N_{1}}\Phi\Big{[}\frac{\alpha^{(i)}+\beta_{new}^{(i)}t_{k}-D}{\sigma_{\epsilon}^{(i)}}\Big{]}}.$
The approximated pdf of $T_{new}|\bm{Y}=\bm{y},T_{new}>0$, is given by :
$f_{T_{new}|\bm{Y}=\bm{y},T_{new}>0}^{approx}(t)=\frac{\frac{1}{N_{1}}\sum_{i=1}^{N_{1}}\phi\Big{[}\frac{\alpha^{(i)}+\beta_{new}^{(i)}(t+t_{k})-D}{\sigma_{\epsilon}^{(i)}}\Big{]}\times\frac{\beta_{new}^{(i)}}{\sigma_{\epsilon}^{(i)}}}{1-\frac{1}{N_{1}}\sum_{i=1}^{N_{1}}\Phi\Big{[}\frac{\alpha^{(i)}+\beta_{new}^{(i)}t_{k}-D}{\sigma_{\epsilon}^{(i)}}\Big{]}}.$ (14)
where $\phi(\cdot)$is the pdf of standard normal random variable.
Observe that if $\beta_{new}^{(i)}>0$ $\forall$ $i$, then $F_{T_{new}|\bm{Y}=\bm{y}}^{approx}(t)$, $F_{T_{new}|\bm{Y}=\bm{y},T_{new}>0}^{approx}(t)$ is a cumulative distribution function and $f_{T_{new}|\bm{Y}=\bm{y},T_{new}>0}^{approx}(t)$ is a probability density function. Observe that the approximated pdf of residual lifetime depends on the samples from the posterior distribution of $\alpha$, $\beta_{new}$ and $\sigma_{\epsilon}^{2}$. In this procedure entire observed degradation measures are utilized to derive residual lifetime distribution of a new unit. The proposed method has an advantage in the sense that is as soon as new degradation signals are available, we update the posterior distribution required parameters and sample from the updated posterior distribution and use the samples to estimate residual lifetime distribution. Next we consider how to generate samples from posterior distribution of required parameters using the proposed Bayesian semi-parametric degradation model.
4 Generating observations from posterior distribution
Here we consider simulation from the posterior distribution. We assume that $\alpha\sim\mathcal{N}(\mu_{\alpha},\sigma_{\alpha}^{2})$ and $\mu_{\alpha}\sim\mathcal{N}(\mu_{1},\sigma_{1}^{2})$, $(\sigma_{\alpha}^{2})^{-1}\sim\mathcal{G}(a_{1},a_{2})$, where we choose $(\mu_{1},\sigma_{1}^{2})=(0,10^{3})$ and $(a_{1},a_{2})=(0.001,0.001)$. It is considered that $\mu_{h}\sim\mathcal{N}(m_{\mu},\sigma_{z})$, $({\sigma_{z}^{2}})^{-1}\sim\mathcal{G}(\tau_{1},\tau_{2})$ , $({\sigma_{h}^{2}})^{-1}\sim\mathcal{G}(a,b)$, $h=1,.....,N$. Available prior information about the degradation rate can be used for choosing hyperparameter $m_{\mu}$. Santos and Loschi [26] used observed data for choosing the values of hyperparameters of $\mu_{h}$ and $\sigma_{h}^{2}$. Motivated by their ideas, we choose $m_{\mu}$, $a$ and $b$ as follows. We fit a simple linear regression model using the expression of equation (2) for the degradation path of each unit. Assuming that there are enough observations to estimate $\beta_{i}$ for each $i$, let $\beta_{i}^{*}$ be the its least square estimate of for $i$th unit, $i=1,...,(n+1)$. Now we consider $m_{\mu}$ and $v_{\beta}$ as the sample average and variance of the $(n+1)$ many $\beta_{i}^{*}$. For the hyper-parameters of prior distribution of $\sigma_{h}^{2}$, we consider $a=\sqrt{v_{\beta}},b={\sqrt[3]{v_{\beta}}}$. For the hyperpriors for $({\sigma_{z}^{2}})^{-1}$, we choose $(\tau_{1},\tau_{2})=(0.001,0.001)$ and for $({\sigma_{\epsilon}^{2}})^{-1}$, we consider $(a_{\epsilon},b_{\epsilon})=(0.001,0.001)$. For the Dirichlet mass parameter $\gamma$, it is assumed that $\gamma\sim G(\eta_{1},\eta_{2})$, where $\eta_{1}=2,\eta_{2}=2$, since this is a good choice as mentioned in Ishwaran [31]. Furthermore it is also assumed that prior distributions are independent.
4.1 Conditional distributions for Gibbs sampling
We generate observations from posterior distribution by Gibbs sampling method. Due to Dirichlet process prior there are ties among the classification variables. Let us consider that ${{K_{1}}^{*},.....,{K_{m}}^{*}}$ denotes the present $m\leq(n+1)$ unique values of K, where $\textbf{K}=(K_{1},.....,K_{n+1})$. In this article, we denote a vector $\textbf{Q}_{(-i)}$ which contains all $Q_{j}$’s where $j\neq i$. In each iteration of the Gibbs sampler we simulate from the following conditional distributions :
  1. 1.
    Conditional distribution of $\alpha$ given $(\bm{\beta},\mu_{\alpha},\sigma_{\alpha}^{2},\bm{\mu},\sigma_{z}^{2},\textbf{K},\textbf{p},\bm{\sigma}^{2},\sigma_{\epsilon}^{2},\gamma,\bm{Y})$ is
    $\mathcal{N}\Big{(}\frac{\sigma_{\alpha}^{2}\sum\limits_{i=1}^{(n+1)}\sum\limits_{j=1}^{n_{i}}(y_{ij}-\beta_{i}t_{ij})+\sigma_{\epsilon}^{2}\mu_{\alpha}}{\sigma_{\alpha}^{2}\sum\limits_{i=1}^{n+1}n_{i}+\sigma_{\epsilon}^{2}},\frac{\sigma_{\alpha}^{2}\sigma_{\epsilon}^{2}}{\sigma_{\alpha}^{2}\sum\limits_{i=1}^{n+1}n_{i}+\sigma_{\epsilon}^{2}}\Big{)}$.
  2. 2.
    Conditional distribution of $\mu_{\alpha}$ given $(\bm{\beta},\alpha,\sigma_{\alpha}^{2},\bm{\mu},\sigma_{z}^{2},\textbf{K},\textbf{p},\bm{\sigma}^{2},\sigma_{\epsilon}^{2},\gamma,\bm{Y})$ is
    $\mathcal{N}\Big{(}\frac{\alpha\sigma_{1}^{2}+\mu_{1}\sigma_{\alpha}^{2}}{\sigma_{1}^{2}+\sigma_{\alpha}^{2}},\frac{\sigma_{\alpha}^{2}\sigma_{1}^{2}}{\sigma_{\alpha}^{2}+\sigma_{1}^{2}}\Big{)}$.
  3. 3.
    Conditional distribution of $(\sigma_{\alpha}^{2})^{-1}$ given $(\bm{\beta},\alpha,\mu_{\alpha},\bm{\mu},\sigma_{z}^{2},\textbf{K},\textbf{p},\bm{\sigma}^{2},\sigma_{\epsilon}^{2},\gamma,\bm{Y})$ is
    $\mathcal{G}\Big{(}a_{1}+\frac{1}{2},b_{1}+\frac{(\alpha-\mu_{\alpha})^{2}}{2}\Big{)}$.
  4. 4.
    Conditional distribution of $\beta_{i}$ given $(\bm{\beta}_{(-i)},\alpha,\mu_{\alpha},\sigma_{\alpha}^{2},\bm{\mu},\sigma_{z}^{2},\textbf{K},\textbf{p},\bm{\sigma}^{2},\sigma_{\epsilon}^{2},\gamma,\bm{Y})$ is
    $\mathcal{N}\Big{(}\frac{\sigma_{K_{i}}^{2}\sum\limits_{j=1}^{n_{i}}t_{ij}(y_{ij}-\alpha)+\sigma_{\epsilon}^{2}\mu_{K_{i}}}{\sigma_{K_{i}}^{2}\sum\limits_{j=1}^{n_{i}}t_{ij}^{2}+\sigma_{\epsilon}^{2}},\frac{\sigma_{K_{i}}^{2}\sigma_{\epsilon}^{2}}{\sigma_{K_{i}}^{2}\sum\limits_{j=1}^{n_{i}}t_{ij}^{2}+\sigma_{\epsilon}^{2}}\Big{)},\hskip 14.22636pti=1,...,n+1$.
  5. 5.
    Conditional distribution of $(\sigma_{\epsilon}^{2})^{-1}$ given $(\bm{\beta},\alpha,\sigma_{\alpha}^{2},\bm{\mu},\sigma_{z}^{2},\textbf{K},\textbf{p},\bm{\sigma}^{2},\gamma,\bm{Y})$ is,
    $\mathcal{G}\Big{(}a_{\epsilon}+\frac{\sum\limits_{i=1}^{(n+1)}n_{i}}{2},b_{\epsilon}+\frac{\sum\limits_{i=1}^{(n+1)}\sum\limits_{j=1}^{n_{i}}(y_{ij}-\alpha-\beta_{i}t_{ij})^{2}}{2}\Big{)}$.
  6. 6.
    Conditional distribution of $K_{i}$ given $(\bm{\beta},\alpha,\sigma_{\alpha}^{2},\bm{\mu},\sigma_{z}^{2},\textbf{K}_{(-i)},\textbf{p},\bm{\sigma}^{2},\sigma_{\epsilon}^{2},\gamma,\bm{Y})$ is,
    $\sum\limits_{h=1}^{N}p_{h,i}\delta_{k}(\cdot)$,       $i=1,...,n+1$,
    where $(p_{1,i},......,p_{N,i})\propto(\frac{p_{1}}{\sigma_{1}}\exp(-\frac{(\beta_{i}-\mu_{1})^{2}}{2\sigma_{1}^{2}}),.....,\frac{p_{N}}{\sigma_{N}}\exp(-\frac{(\beta_{i}-\mu_{N})^{2}}{2\sigma_{N}^{2}}))$.
  7. 7.
    Conditional distribution of p given $(\bm{\beta},\mu_{\alpha},\sigma_{\alpha}^{2},\bm{\mu},\sigma_{z}^{2},\textbf{K},\bm{\sigma}^{2},\sigma_{\epsilon}^{2},\gamma,\bm{Y})$ is represented in terms of distribution of $V_{h}^{*}$’s, $h=1,...,(N-1)$, where $p_{1}=V_{1}^{*}$ and $p_{h}=(1-V_{1}^{*})(1-V_{2}^{*})...(1-V_{h-1}^{*})V_{h}^{*}\hskip 2.84526pt,h=2,...,(N-1)$ and $p_{N}=1-\sum\limits_{h=1}^{N-1}p_{h}$ and the $V_{h}^{*}$ are independent and $V_{h}^{*}\sim Beta(1+r_{h},\gamma+\sum\limits_{l=k+1}^{N}r_{l})$, where $r_{h}=\sum\limits_{i=1}^{n}I_{K_{i}=h}$, $h=1,.......,N$.
  8. 8.
    Conditional distribution of ${\mu}_{h}$ given $(\bm{\beta},\alpha,\sigma_{\alpha}^{2},\bm{\mu}_{(-h)},\sigma_{z}^{2},\textbf{K},\textbf{p},\bm{\sigma}^{2},\sigma_{\epsilon}^{2},\gamma,\bm{Y}=\bm{y})$, for each $h\in[{K_{1}}^{*},.....,{K_{m}}^{*}]$ is $N(\mu_{h}^{*},{\sigma_{h}^{2}}^{*})$, where $\mu_{h}^{*}={\sigma_{h}^{2}}^{*}\times\sum\limits_{i:K_{i}=h}(X_{i}/{\sigma_{h}^{2}}+m_{\mu}/\sigma_{z})$, and ${\sigma_{h}^{2}}^{*}=\frac{{\sigma_{h}^{2}}{\sigma_{z}^{2}}}{r_{h}{\sigma_{z}^{2}}+{\sigma_{h}^{2}}}$. For $h\in\textbf{K}-[{K_{1}}^{*},.....,{K_{m}}^{*}]$ independently simulate from $N(m_{\mu},{\sigma_{z}^{2}})$.
  9. 9.
    Conditional distribution of $(\sigma_{z}^{2})^{-1}$ given $(\bm{\beta},\mu_{\alpha},\sigma_{\alpha}^{2},\bm{\mu},\textbf{K},\textbf{p},\bm{\sigma}^{2},\sigma_{\epsilon}^{2},\gamma,\bm{Y})$ is,
    $\mathcal{G}(\tau_{1}+\frac{N}{2},\tau_{2}+\sum\limits_{h=1}^{N}(\mu_{h}-m_{\mu})^{2})$.
  10. 10.
    Conditional distribution of $(\sigma_{h}^{2})^{-1}$ given $(\bm{\beta},\alpha,\sigma_{\alpha}^{2},\bm{\mu},\sigma_{z}^{2},\textbf{K},\textbf{p},\bm{\sigma}^{2}_{(-h)},\sigma_{\epsilon}^{2},\gamma,\bm{Y})$, for each $h\in[{K_{1}}^{*},.....,{K_{m}}^{*}]$ is $\mathcal{G}(a+\frac{r_{h}}{2},b+\sum\limits_{i:K_{i}=h}(\beta_{i}-\mu_{h})^{2})$. For $h\in\textbf{K}-[{K_{1}}^{*},.....,{K_{m}}^{*}]$ independently simulate from $\mathcal{G}(a,b)$.
  11. 11.
    Conditional distribution of concentration parameter $\gamma$ given $(\bm{\beta},\mu_{\alpha},\sigma_{\alpha}^{2},\bm{\mu},\sigma_{z}^{2},\textbf{K},\textbf{p},\bm{\sigma}^{2},\sigma_{\epsilon}^{2},\bm{Y})$ is,
    $\mathcal{G}(N+\eta_{1}-1,\eta_{2}-\sum\limits_{k=1}^{N-1}\log(1-V_{k}^{*}))$.
We simulate total 200000 MCMC samples from the posterior distributions and discard initial 50000 samples as burn-in. We consider only the observations with lag sizes 50, to decrease the autocorrelation. The convergence of the generated Markov chains is validated by auto-correlation plot and trace plot. We compute the residual lifetime distribution of a unit based on the generated observations from posterior distribution of parameters. Next we discuss on generating observations from the derived residual lifetime distribution.
4.2 Transformation based Markov chain Monte Carlo (TMCMC)
Once we estimate the probability density function of residual lifetime distribution, our next goal is to predict residual lifetime for a unit functioning at some point $t_{k}$. One way to predict is simulating observation from the estimated distribution and a specific sample quantile of that generated samples can be used as predicted residual lifetime. MCMC techniques like Metropolis-Hastings (MH) helps to simulate observation from a given target distribution. However, several challenges arise for implementation of this algorithm. For example, a very large number of iterations (of the order of millions) are usually necessary for simulating observations from the target distribution. Another key issue for implementing MH algorithm is that it often leads to poor acceptance rates. To overcome these issues, we use a MCMC technique called Transformation based Markov chain Monte Carlo (TMCMC) algorithm introduced by Dutta and Bhattacharya [30] for simulating samples from the estimated residual lifetime distribution. This algorithm produces a rapidly mixing Markov chain and a better acceptance rate compared to the standard MH algorithms. For univariate case, TMCMC can be reduced to a MH algorithm with a specific proposal distribution, though in higher dimensions the proposal does not admit a mixture form and TMCMC cannot be a special case of the MH algorithm.
Consider $S$ be the state space and suppose $T:S\times S^{\prime}\rightarrow S$ for some $S^{\prime}$ (may be a subset of $S$) is a differentiable transformation. TMCMC technique is based on constructing forward and backward transformations $T$ which are to be defined in such a way that the detailed balance and irreducibility hold for the Markov chain generated by this technique. If the current state is $x$, then the forward move is proposed by $x^{\prime}=x\epsilon$, where $\epsilon\in(0,1)$ is a simulated value from some arbitrary density of the form $g(\epsilon)I_{\epsilon}(0,1)$. The backward transformation $\frac{x^{\prime}}{\epsilon}$ is applied for moving back to $x$ from $x^{\prime}$. In general, for given $\epsilon$ and the current state $x$, forward transformation is denoted by $T(x,\epsilon)$, and the backward transformation by $T^{b}(x,\epsilon)$. One can observe that, the regions covered by the two transformations are disjoint. An important advantage associated with this algorithm is that whatever be the choice of the density $g(\epsilon)$, it cancels in the acceptance ratio of the proposed TMCMC algorithm. For more details see Dutta and Bhattacharya [30].
The transformation $T$ is considered to be a differentiable transformation and the corresponding Jacobian is constructed as $J(x,\epsilon)=|\frac{\partial(T(x,\epsilon),\epsilon)}{\partial(x,\epsilon)}|$ which is non-zero almost everywhere. In this article we use the transformation, where $S=S^{\prime}=(0,\infty)$ and $T(x,\epsilon)=x\epsilon$. For all $x\in S$, $T^{b}(x,\epsilon)=\frac{x}{\epsilon}$; $A=(0,1)$. In this case Jacobian is $\epsilon$. Suppose $g$ is a density on $A$ where $A$ is a subset of $S^{\prime}$ such that $T(x,A)$ and $T^{b}(x,A)$ are disjoint and $10<p<$ 1, $\alpha(\cdot)$ be the acceptance ratio and $\pi(\cdot)$ be the target distribution. Consider an initial value $x_{0}$ and $x_{t}$ be the value that the chain takes at $t$th iteration, where $t\geq 1$. Then the MCMC algorithm based on transformation is constructed as follows.
  1. 1.
    Generate $\epsilon\sim g(\cdot)$ and $u\sim U(0,1)$ independently.
  2. 2.
    if $u<p$, $x^{\prime}=T(x_{t},\epsilon)$ and $\alpha(x_{t},\epsilon)=min(1,\frac{1-p}{p}\frac{\pi(x^{\prime})}{\pi(x)}J(x,\epsilon))$
    else if $p<u<1$, $x^{\prime}=T^{b}(x_{t},\epsilon)$ and $\alpha(x_{t},\epsilon)=min(1,\frac{p}{1-p}\frac{\pi(x^{\prime})}{\pi(x)}\frac{1}{J(x,\epsilon)})$
  3. 3.
    set, $x_{t+1}=\left\{\begin{array}[]{ll}x^{\prime},\hskip 5.69054pt\text{with probability}\hskip 5.69054pt\alpha(x_{t},\epsilon).\\ x_{t},\hskip 5.69054pt\text{with probability}\hskip 5.69054pt1-\alpha(x_{t},\epsilon).\end{array}\right.$
  4. 4.
    Repeat Steps 1-3 $N_{2}$ times.
In this study the target probability density function is the density function as presented in equation (14). We have simulated 50000 MCMC iterates after discarding the initial 5000 iterations as burn-in. We take samples from lag size 10 from the rest of the samples and use the median of this as our predicted value of residual lifetime. Consider the predicted life as ${T_{ip}}$ and actual life as $T_{ia}$ for the $i$th individual unit. The prediction accuracy for $i$th individual unit is assessed by the relative absolute error $\frac{|T_{ip}-T_{ia}|}{T_{ia}}$. Figure 1 represents framework of our proposed Bayesian semi-parametric methodology which comprises of modeling degradation data to predicting residual lifetime of an individual unit.
Degradation data of $n$ units and a new unit with some degradation observations General Path model using $(n+1)$ units is built ( A mixed effect model) Dirichlet process mixture prior for the random effect distribution and parametric priors for fixed effect and variance of measurement errors is considered Simulate from the posterior distribution of parameters using Gibbs sampling as described above in Section 4.1 Use the posterior samples to compute the residual lifetime distribution of the new unit Use TMCMC algorithm to simulate observations from a derived distribution Compute median of the samples generated and use this as predicted value of residual lifetime and compute accuracy of prediction
Figure 1: Framework of proposed method
5 Simulation study
In this section we perform two simulation experiments to demonstrate the performance of the proposed Bayesian semi-parametric degradation modeling approach. We consider two and three component mixture distributions for random effects for two experiments. For the first case, we consider the true random effect distribution as $\beta\sim 0.6N(3,0.1)+0.4N(6,0.2)$. We consider true fixed effect $\alpha$ as 2 and the distribution of error represented as $\epsilon_{ij}\overset{\mathrm{i.i.d}}{\sim}N(0,0.5)$. The threshold value $D$ ts taken as 9. We generate degradation measures for 10 units and for each unit the degradation is measured in the time interval $(0,1)$ at equally spaced times $\frac{1}{10}$, until the degradation measure crosses the threshold value $D=9$, that is if for some unit degradation measures crosses threshold before the maximum time which is 1 in this case, we stop observing data and which implies that $n_{i}\leq 10$ for $i=1,...,10$. We assume the true linear model $Y_{ij}=\alpha+\beta_{i}t_{ij}+\epsilon_{ij},\hskip 5.69054pti=1,....,10,j=1,....,n_{i}$. For the second case we consider the true random effect distribution with mixture of 3 normal distribution represented as $\beta\sim 0.4N(2,0.1)+0.3N(3,0.15)+0.3N(6,0.12)$. Here also we take fixed effect $\alpha$ as 2 and $\epsilon_{ij}\overset{\mathrm{i.i.d}}{\sim}N(0,0.6)$ and $D$ as 9. We generate degradation measures for 30 units and for each unit the degradation is measured according to the plan followed for case 1.
We run the simulation procedure to generate degradation path from the above-mentioned models in both the cases. We generate observations using the sample interval of 0.001 and obtain the time when the path crosses the predefined threshold for the first time. In this work we predict residual lifetime of each unit at the last time point when the observed degradation measurement has not crossed threshold value. We evaluate the prediction accuracy of the residual life using the error criteria introduced in the Section 4.2. The proposed Bayesian semi-parametric method is compared with a Bayesian parametric method. For Bayesian parametric method, we consider $\beta_{i}\sim N(\mu_{\beta},\sigma_{\beta}^{2})$, and $\alpha\sim N(\mu_{\alpha},\sigma_{\alpha}^{2})$ and $(\sigma_{\epsilon}^{2})^{-1}\sim G(a_{\epsilon},b_{\epsilon})$. The prior specification is: $\mu_{\beta}\sim N(m_{\mu},\sigma_{z}^{2})$, $({\sigma_{z}^{2}})^{-1}\sim\mathcal{G}(\tau_{1},\tau_{2})$ and $(\sigma_{\beta}^{2})^{-1}\sim G(a,b)$ where $a=\sqrt{v_{\beta}},b=\sqrt[3]{v_{\beta}}$. We choose $\mu_{\alpha}\sim N(0,10^{3})$ and $(\sigma_{\alpha}^{2})^{-1}\sim G(0.001,0.001)$. We implement a similar procedure discussed in section 4 to choose values for $m_{\mu}$ and $v_{\beta}$. $(\tau_{1},\tau_{2})=(0.001,0.001)$ and $(a_{\epsilon},b_{\epsilon})=(0.001,0.001)$ is considered for this case. The degradation path simulated from the two cases are represented in Figure 2.
Figure 2: Degradation paths generated from two different cases: (a) Case 1 (b) Case 2.
5.1 Case 1
For the first experiment we consider a 2-component mixture of normal distribution for random effect where mean of one component is twice the mean of the other component. So we expect that some components degrade significantly faster than others, which is also evident from Figure 2(a). For semi-parametric method we take truncation number $N$ as the number of units in the study that is 10. The mean, standard deviation (s.d) and 95% Credible interval of random effects for each of 10 units, fixed effect $\alpha$ and variance of error $\sigma_{\epsilon}^{2}$ computed based on generated posterior samples are presented in Table 1.
Table 1: Posterior mean, standard deviation and 95% Credible Interval of random effects ($\beta_{i}$, $i=1,...,10$ ), fixed effect ($\alpha$) and variance of measurement error ($\sigma_{\epsilon}^{2}$)
Parameter Parametric method Semi-parametric method
mean s.d 95% Credible interval mean s.d 95% Credible interval
$\beta_{1}$ 3.048 0.465 (2.137, 3.981) 2.968 0.458 (2.054, 3.864)
$\beta_{2}$ 5.485 0.463 (4.571, 6.410) 5.238 0.479 (4.289, 6.198)
$\beta_{3}$ 2.309 0.466 (1.397, 3.246) 2.285 0.487 (1.345, 3.278)
$\beta_{4}$ 3.559 0.451 (2.690, 4.458) 3.453 0.452 (2.590, 4.346)
$\beta_{5}$ 4.192 0.449 (3.304, 5.075) 4.041 0.458 (3.166, 4.927)
$\beta_{6}$ 5.627 0.457 (4.763, 6.529) 5.341 0.488 (4.391, 6.293)
$\beta_{7}$ 2.955 0.465 (2.060, 3.910) 2.903 0.457 (2.016, 3.771)
$\beta_{8}$ 3.324 0.455 (2.469, 4.260) 3.216 0.454 (2.312, 4.103)
$\beta_{9}$ 3.365 0.453 (2.478, 4.263) 3.246 0.455 (2.345, 4.118)
$\beta_{10}$ 6.960 0.472 (6.033, 7.928) 6.596 0.538 (5.521, 7.616)
$\alpha$ 1.878 0.141 (1.591, 2.154) 1.956 0.154 (1.662, 2.261)
$\sigma_{\epsilon}^{2}$ 0.707 0.141 (0.481, 1.037) 0.677 0.142 (0.443, 0.986)
It is evident from Table 1, that for the proposed semi-parametric method the estimated posterior mean of fixed effect is closer to the true value compared to the parametric method. On the other side, estimated posterior mean for error variance is higher than true value for both the methods. It is observed that the units that are slowly degrading is characterized by random effects with low posterior means and significantly higher posterior means corroborate to the fact that the unit is degrading faster. Observe that for fast degrading units, posterior mean of random effect is little higher for parametric method compared to semi-parametric method. Also for the slowly degrading units the same pattern follows although the posterior mean is very close for both the methods. The posterior samples generated are used to estimate residual lifetime distribution of each individual unit. For each unit we apply the above-mentioned TMCMC technique to simulate observations from the target distribution. For each case we take $p=\frac{1}{2}$.
Figure 3: For each row, the left side plot is for histogram for Residual lifetimes, middle one is for trace plot and right side denotes autocorrelation plot of the samples produced by semi-parametric method.
Figure 4: For each row, the left side plot is for histogram for Residual lifetimes, middle one is for trace plot and right side denotes autocorrelation plot of the samples produced by parametric method.
Table 2: Estimated median and 95% predictive interval of residual lifetime using a Bayesian parametric method and proposed Bayesian semi-parametric method
Unit True value Parametric method Semi-parametric method
median 95% Predictive interval median 95% Predictive interval
Unit 1 0.665 1.362 (0.548, 2.388) 1.417 (0.514, 2.490)
Unit 2 0.023 0.211 (0.001, 0.428) 0.222 (0.001, 0.453)
Unit 3 0.894 2.094 (0.812, 3.967) 2.128 (0.789, 4.238)
Unit 4 0.710 1.012 (0.378, 1.812) 1.055 (0.385, 1.855)
Unit 5 0.607 0.700 (0.207, 1.309) 0.745 (0.248, 1.380)
Unit 6 0.024 0.175 (0.001, 0.377) 0.209 (0.001, 0.441)
Unit 7 0.886 1.457 (0.618, 2.599) 1.437 (0.520, 2.545)
Unit 8 0.917 1.158 (0.449, 2.028) 1.191 (0.407, 2.125)
Unit 9 0.551 1.122 (0.413, 1.968) 1.164 (0.521, 2.153)
Unit 10 0.025 0.158 (0.001, 0.393) 0.162 (0.001, 0.394)
The samples produced by TMCMC technique for both parametric and semi-parametric model are used for prediction. We collect the samples according to the discussion done in Section 4.2 and they are used for prediction. Predictive intervals are constructed using the 95% highest probability density region of the generated samples. The estimates for residual lifetime are rounded off to three decimal places are presented in Table 2 along with the true residual lifetime values for each unit. Figure 3 and 4 shows histograms of the generated residual lifetimes of Unit 1,2 and 3 along with the trace plot of and autocorrelation plot of generated samples for semi-parametric and parametric method respectively. Histograms and autocorrelation plots are generated by samples at lag size 10 for each unit. In each case we consider the median of the samples as the predicted residual lifetime value for each individual unit. It is evident from Table 2 that for most of the units both parametric and semi-parametric method produce intervals containing the true value.
We observe that error in predicting the residual lifetime using both the proposed semi-parametric and parametric method is nearly equal for slowly degrading units. On the other side, prediction error for parametric method is lesser than semi-parametric method for faster degrading units. Comparison of prediction error for both the methods is provided in Appendix D.
5.2 Case 2
In second case we consider a 3-component mixture of normal distribution for random effect, so we expect three different cluster of units with significant different degradation rates, which is also evident from Figure 2(b). For semi-parametric method we take truncation number $N$ as 30. The mean, standard deviation (s.d) and 95% credible interval computed based on generated posterior samples for random effects, fixed effect $\alpha$ and variance of measurement error $\sigma_{\epsilon}^{2}$ are presented in Table 3.
Table 3: Posterior mean, standard deviation and 95% Credible Interval of random effects ($\beta_{i}$, $i=1,...,30$), fixed effect ($\alpha$) and variance of measurement error ($\sigma_{\epsilon}^{2}$)
Parameter Parametric method Semi-parametric method
mean s.d 95% Credible interval mean s.d 95% Credible interval
$\beta_{1}$ 4.374 0.453 (3.496, 5.270) 4.388 0.435 (3.553, 5.269)
$\beta_{2}$ 3.723 0.443 (2.884, 4.611) 3.808 0.450 (2.946, 4.712)
$\beta_{3}$ 2.408 0.439 (1.531, 3.250) 2.642 0.491 (1.712, 3.622)
$\beta_{4}$ 4.034 0.441 (3.146, 4.874) 4.096 0.446 (3.237, 5.001)
$\beta_{5}$ 2.811 0.441 (1.946, 3.683) 3.003 0.463 (2.145, 3.941)
$\beta_{6}$ 5.429 0.430 (4.589, 6.281) 5.330 0.468 (4.367, 6.130)
$\beta_{7}$ 3.373 0.444 (2.519, 4.242) 3.511 0.456 (2.653, 4.405)
$\beta_{8}$ 2.586 0.443 (1.744, 3.494) 2.804 0.477 (1.881, 3.805)
$\beta_{9}$ 4.886 0.460 (3.977, 5.813) 4.834 0.447 (3.941, 5.711)
$\beta_{10}$ 5.316 0.442 (4.485, 6.218) 5.240 0.458 (4.393, 6.149)
$\beta_{11}$ 2.627 0.449 (1.750, 3.510) 2.843 0.490 (1.909, 3.826)
$\beta_{12}$ 6.197 0.445 (5.351, 7.072) 6.017 0.504 (5.002, 7.011)
$\beta_{13}$ 4.743 0.434 (3.897, 5.584) 4.703 0.440 (3.849, 5.593)
$\beta_{14}$ 2.529 0.450 (1.656, 3.412) 2.781 0.486 (1.817, 3.757)
$\beta_{15}$ 4.127 0.443 (3.288, 5.014) 4.180 0.443 (3.304, 5.068)
$\beta_{16}$ 1.337 0.453 (0.465, 2.247) 1.696 0.559 (0.684, 2.861)
$\beta_{17}$ 2.320 0.445 (1.470, 3.225) 2.572 0.494 (1.632, 3.553)
$\beta_{18}$ 3.554 0.443 (2.670, 4.415) 3.674 0.449 (2.803, 4.579)
$\beta_{19}$ 4.023 0.433 (3.177, 4.887) 4.083 0.449 (3.194, 4.985)
$\beta_{20}$ 2.187 0.438 (1.315, 3.058) 2.453 0.506 (1.475, 3.470)
$\beta_{21}$ 3.408 0.437 (2.549, 4.278) 3.551 0.463 (2.667, 4.501)
$\beta_{22}$ 4.769 0.433 (3.921, 5.644) 4.755 0.458 (3.885, 5.674)
$\beta_{23}$ 2.920 0.442 (2.055, 3.799) 3.097 0.471 (2.177, 4.022)
$\beta_{24}$ 4.325 0.440 (3.492, 5.165) 4.342 0.447 (3.470, 5.237)
$\beta_{25}$ 2.191 0.439 (1.328, 3.063) 2.454 0.494 (1.499, 3.438)
$\beta_{26}$ 6.698 0.448 (5.833, 7.579) 6.469 0.518 (5.445, 7.468)
$\beta_{27}$ 5.920 0.442 (5.062, 6.786) 5.772 0.483 (4.852, 6.735)
$\beta_{28}$ 4.486 0.440 (3.639, 5.362) 4.505 0.438 (3.674, 5.404)
$\beta_{29}$ 1.553 0.446 (0.677, 2.453) 1.872 0.548 (0.834, 2.999)
$\beta_{30}$ 6.633 0.441 (5.786, 7.508) 6.391 0.516 (5.392, 7.419)
$\alpha$ 2.091 0.086 (1.922, 2.257) 2.051 0.116 (1.813, 2.268)
$\sigma_{\epsilon}^{2}$ 0.745 0.079 (0.605, 0.921) 0.761 0.074 (0.577, 1.028)
It is evident from Table 3 that for the proposed semi-parametric method the estimated posterior mean of fixed effect is closer to the true value for both parametric and semi-parametric method. The estimated posterior mean for error variance is higher than true value for both the methods, which is also evident in the Case 1 analysis. Observe that for fastest degrading units, posterior mean of random effect is higher for parametric method compared to semi-parametric method, for example see units 6 and 27. For unit 2 and 4 which can be considered as a moderate speed degrading units, it is evident that the semi-parametric produces higher posterior means of random effects compared to parametric model. The same pattern is also evident for the slow degrading units also, for example see units 20 and 25.
Figure 5: For each row, the left side plot is for histogram for Residual lifetimes, middle one is for trace plot of generated samples and right side denotes autocorrelation plot produced by Semi-parametric method.
Figure 6: For each row, the left side plot is for histogram for Residual lifetimes, middle one is for trace plot of generated samples and right side denotes autocorrelation plot produced by parametric method.
Table 4: Estimated median and 95% predictive interval of residual lifetime using a Bayesian parametric method and proposed Bayesian semi-parametric method
Unit True value Parametric method Semi-parametric method
median 95% Predictive interval median 95% Predictive interval
Unit 1 0.071 0.579 (0.127, 1.087) 0.487 (0.094, 0.924)
Unit 2 0.259 0.841 (0.265, 1.514) 0.675 (0.196, 1.220)
Unit 3 1.261 1.832 (0.735, 3.372) 1.626 (0.635, 2.935)
Unit 4 0.354 0.711 (0.185, 1.310) 0.688 (0.155, 1.237)
Unit 5 0.508 1.472 (0.601, 2.615) 1.089 (0.392, 1.958)
Unit 6 0.021 0.286 (0.001, 0.624) 0.321 (0.002, 0.680)
Unit 7 0.858 1.035 (0.308, 1.827) 0.978 (0.341, 1.768)
Unit 8 0.842 1.651 (0.613, 2.958) 1.454 (0.524, 2.719)
Unit 9 0.021 0.158 (0.020, 0.322) 0.166 (0.016, 0.326)
Unit 10 0.023 0.210 (0.003, 0.438) 0.226 (0.001, 0.471)
Unit 11 0.466 1.638 (0.657, 3.028) 0.959 (0.334, 1.724)
Unit 12 0.022 0.231 (0.001, 0.502) 0.267 (0.002, 0.562)
Unit 13 0.038 0.228 (0.034, 0.461) 0.241 (0.033, 0.471)
Unit 14 0.984 1.742 (0.727, 3.175) 1.511 (0.563, 2.800)
Unit 15 0.094 0.334 (0.086, 0.611) 0.335 (0.066, 0.612)
Unit 16 1.968 4.243 (1.532, 10.864) 3.166 (0.855, 7.597)
Unit 17 1.060 1.981 (0.787, 3.591) 1.684 (0.616, 3.103)
Unit 18 0.152 0.966 (0.327, 1.703) 0.348 (0.188, 0.998)
Unit 19 0.096 0.288 (0.074, 0.526) 0.282 (0.072, 0.516)
Unit 20 1.398 2.133 (0.820, 3.991) 1.827 (0.598, 3.436)
Unit 21 0.149 1.034 (0.407, 1.909) 0.596 (0.222, 1.110)
Unit 22 0.024 0.187 (0.008, 0.358) 0.193 (0.016, 0.377)
Unit 23 0.339 1.370 (0.556, 2.407) 0.780 (0.271, 1.409)
Unit 24 0.088 0.236 (0.052, 0.452) 0.243 (0.036, 0.451)
Unit 25 0.994 2.139 (0.800, 3.975) 1.813 (0.675, 3.473)
Unit 26 0.022 0.167 (0.001, 0.419) 0.199 (0.001, 0.475)
Unit 27 0.021 0.199 (0.001, 0.488) 0.226 (0.001, 0.541)
Unit 28 0.036 0.227 (0.045, 0.442) 0.226 (0.038, 0.433)
Unit 29 1.037 3.458 (1.253, 7.828) 2.721 (0.813, 6.098)
Unit 30 0.021 0.116 (0.001, 0.328) 0.148 (0.001, 0.423)
Similar to the previous analysis in Case 1, we generate observations from the residual lifetime distribution using TMCMC technique with $p=\frac{1}{2}$. Predictive intervals are constructed for each unit using the 95% highest probability density region of the generated samples. As in the previous case, we consider the median of the samples as the predicted value for each individual unit. Table 4 presents residual lifetime estimates rounded off to three decimal places along with the true values for each 30 unit. Histograms of the generated residual lifetimes of Unit 1,2 and 3 along with the trace plot and autocorrelation plot for the generated samples for semi-parametric and parametric method is presented in Figure 5 and 6 respectively. Histograms and autocorrelation plots are generated by samples at lag size 10 for each unit in this case also. We observe that in case of fast degrading units, predicted value by the parametric method is more accurate compared to the semi-parametric method for example see Unit 26,30. For the slow degrading units, it is evident that the predicted residual lifetime by the parametric method is higher than true values compared to the semi-parametric method, for example, see unit 8,11. Similar pattern is also observed for the moderate degrading units, for example see unit 1,4. But there are also some moderate degrading units for example unit 9, 13 where the predicted value for both the method is very close to each other. It is also observed that for most of the cases produced predictive intervals for both the methods contain true value. Comparison of prediction error for the both parametric and semi-parametric method is provided in Appendix D.
6 Data Analysis
Now we demonstrate analysis of fatigue crack size dataset given in Lu & Meeker[2]. The measurements are taken for 21 units. For each unit the initial crack size was 0.9 inch and data collection was terminated at the first inspection after a unit’s crack size reached 1.6 inches or censored after 0.12 million cycles, whichever came first. Degradation paths of the units are presented in Figure 7. It may be noted from the plot that some of the units are degrading faster than other units, so it can be considered that the units are coming from heterogeneous population. For semi-parametric method we take truncation number $N$ as 21.
Figure 7: Development of crack sizes over time [2]
Table 5: Posterior mean, standard deviation and 95% Credible Interval of random effects ($\beta_{i}$, $i=1,...,21$ ), fixed effect ($\alpha$) and variance of measurement error ($\sigma_{\epsilon}^{2}$)
Parameter Parametric method Semi-parametric method
mean s.d 95% Credible interval mean s.d 95% Credible interval
$\beta_{1}$ 5.743 0.784 (4.298, 7.330) 6.211 1.268 (4.179, 8.928)
$\beta_{2}$ 5.649 0.768 (4.255, 7.152) 6.018 1.279 (3.897, 8.845)
$\beta_{3}$ 5.724 0.758 (4.269, 7.289) 6.143 1.141 (4.360, 8.676)
$\beta_{4}$ 5.692 0.780 (4.237, 7.328) 6.088 1.142 (4.203, 8.632)
$\beta_{5}$ 5.671 0.774 (4.146, 7.310) 6.019 1.197 (4.047, 8.581)
$\beta_{6}$ 5.633 0.761 (4.225, 7.219) 5.953 1.154 (3.964, 8.361)
$\beta_{7}$ 5.584 0.754 (4.088, 7.015) 5.924 1.149 (3.989, 8.442)
$\beta_{8}$ 5.561 0.763 (4.075, 7.007) 5.809 1.171 (3.870, 8.409)
$\beta_{9}$ 5.579 0.763 (4.048, 7.118) 5.873 1.113 (3.980, 8.048)
$\beta_{10}$ 5.528 0.743 (4.027, 6.960) 5.770 1.108 (3.910, 8.085)
$\beta_{11}$ 5.415 0.761 (3.930, 6.934) 5.665 1.100 (3.791, 7.968)
$\beta_{12}$ 5.402 0.742 (3.907, 6.872) 5.630 1.147 (3.681, 7.926)
$\beta_{13}$ 5.252 0.749 (3.718, 6.727) 5.404 1.149 (3.467, 7.879)
$\beta_{14}$ 5.212 0.744 (3.698, 6.693) 5.316 1.153 (3.203, 7.660)
$\beta_{15}$ 5.244 0.729 (3.741, 6.537) 5.402 1.149 (3.430, 7.872)
$\beta_{16}$ 5.105 0.745 (3.564, 6.477) 5.194 1.202 (3.099, 7.615)
$\beta_{17}$ 5.051 0.750 (3.632, 6.429) 5.099 1.204 (2.986, 7.531)
$\beta_{18}$ 4.987 0.783 (3.364, 6.405) 5.033 1.179 (2.916, 7.587)
$\beta_{19}$ 4.915 0.800 (3.171, 6.354) 4.942 1.212 ( 2.739, 7.298)
$\beta_{20}$ 4.907 0.777 (3.321, 6.329) 4.870 1.213 (2.661, 7.278)
$\beta_{21}$ 4.853 0.780 (3.208, 6.329) 4.854 1.235 (2.586,7.146)
$\alpha$ 0.844 0.028 (0.787, 0.901) 0.832 0.049 (0.716, 0.899)
$\sigma_{\epsilon}^{2}$ 0.113 0.017 (0.079, 0.149) 0.127 0.048 (0.081, 0.225)
Figure 8: For each row, the left side plot is for histogram for Residual lifetimes, middle one is for trace plot of generated samples and right side denotes autocorrelation plot produced by Semi-parametric method.
Figure 9: For each row, the left side plot is for histogram for Residual lifetimes, middle one is for trace plot of generated samples and right side denotes autocorrelation plot produced by parametric method.
The mean, standard deviation (s.d) and 95% credible interval (C.I) of posterior samples of random effects for each of 21 units, fixed effect $\alpha$ and variance of measurement error $\sigma_{\epsilon}^{2}$ are presented in Table 5. Observe that for some units, posterior mean of random effect is significantly higher than some other units. For example for unit 1, the estimated posterior mean for random effects is much higher than that of unit 20, which can motivate us to analyze this data using the proposed semi-parametric model.
Table 6: Estimated median and 95% predictive interval of residual lifetime using a Bayesian parametric method and proposed Bayesian semi-parametric method
Unit True value Parametric method Semi-parametric method
median 95% Predictive interval median 95% Predictive interval
Unit 1 0.008 0.073 (0.0002, 0.177) 0.041 (0.0002, 0.157)
Unit 2 0.010 0.066 (0.0001, 0.173) 0.035 (0.0001, 0.143)
Unit 3 0.001 0.061 (0.0004, 0.164) 0.020 (0.0001, 0.086)
Unit 4 0.003 0.059 (0.0003, 0.161) 0.025 (0.0001, 0.107)
Unit 5 0.003 0.062 (0.0004, 0.169) 0.016 (0.0001, 0.066)
Unit 6 0.006 0.061 (0.0004, 0.165) 0.019 (0.0001, 0.078)
Unit 7 0.006 0.062 (0.0004, 0.172) 0.055 (0.0002, 0.168)
Unit 8 0.009 0.062 (0.0004, 0.167) 0.033 (0.0001, 0.132)
Unit 9 0.003 0.056 (0.0002, 0.163) 0.015 (0.0001, 0.063)
Unit 10 0.005 0.058 (0.0003, 0.159) 0.014 (0.0001, 0.058)
Unit 11 0.008 0.058 (0.0003, 0.159) 0.021 (0.0001, 0.087)
Unit 12 0.008 0.059 (0.0003, 0.168) 0.025 (0.0001, 0.101)
Unit 13 0.009 0.055 (0.0003, 0.158) 0.059 (0.0004, 0.189)
Unit 14 0.013 0.052 (0.0002, 0.157) 0.027 (0.0001, 0.114)
Unit 15 0.018 0.055 (0.0003, 0.161) 0.035 (0.0001, 0.144)
Unit 16 0.024 0.056 (0.0003, 0.165) 0.060 (0.0003, 0.223)
Unit 17 0.026 0.056 (0.0002, 0.167) 0.064 (0.0004, 0.211)
Unit 18 0.031 0.057 (0.0003, 0.170) 0.066 (0.0003, 0.211)
Unit 19 0.040 0.058 (0.0002, 0.168) 0.070 (0.0004, 0.221)
Unit 20 0.047 0.057 (0.0003, 0.169) 0.072 (0.0003, 0.228)
Unit 21 0.050 0.059 (0.0003, 0.180) 0.072 (0.0004, 0.232)
We derive residual lifetime distribution corresponding to the last time point at which degrading unit has not crossed the threshold value. Predictive intervals are constructed using the 95% highest probability density region of the generated samples using TMCMC method with $p=\frac{1}{2}$. Table 6 presents residual lifetime estimates where the lower limit of the predictive interval is rounded off to four decimal places and rest of the values are rounded off to three decimal places along with the true values for each of 21 units. Histograms of the generated residual lifetimes of first three units along with the trace plot and autocorrelation plot for the generated samples for semi-parametric and parametric method is presented in Figure 8 and 9 respectively. Histograms and autocorrelation plots are constructed based on the samples at lag size 10 for each unit. It is observed that the proposed semi-parametric method is performing better than the parametric method in almost all the cases in terms of predicting residual lifetime of the individual units. It is also evident that for most of the cases the produced predictive intervals for both the methods contain true value. Comparison of prediction error for the both parametric and semi-parametric method is provided in Appendix D.
7 Conclusion
The objective of this article is to predict residual lifetime of units in a heterogeneous situation. To deal with heterogeneity we proposed a Bayesian semi-parametric degradation model. In the first part we assumed a general path model where the random effects are modeled using the Dirichlet process mixture of normal distributions. Model hierarchy is represented according to Ishwaran & Zarrerpur [31]. Gibbs sampling is used to draw samples from the posterior distribution of the model parameters. In the second part of the problem we developed the residual lifetime distribution of each unit which depends on the samples drawn from posterior distribution. Finally we simulate samples from this distribution using MCMC technique and considered the sample median as the predicted residual lifetime of each unit. For implementation of the proposed method we use R programming language. The proposed Bayesian semi-parametric method is compared with a Bayesian parametric method with respect to the error criteria discussed earlier in this paper. In simulation study it is found that, in some of the cases our proposed method is performing well compared to the parametric method whereas there are some cases where parametric method performs better than the proposed method. Finally, we applied the model to Fatigue Crack-Size dataset to evaluate the performance of our model. In this case, it is found that the prediction is more accurate for the proposed semi-parametric method compared to the parametric method for most of the units. The proposed Bayesian semi-parametric degradation model may be used for similar problems like in medical diagnosis contexts as well.
Appendix A Histogram of Residual lifetime, Monitoring MCMC convergence for Bayesian semi-parametric method
A.1 Case 1
Figure 10: For each row, the left side plot is for histogram for Residual lifetimes, middle one is for trace plot and right side denotes autocorrelation plot produced based on generated samples for unit 4-10 by semi-parametric method. Histograms and autocorrelation plots are constructed based on the samples at lag size 10 for each unit.
A.2 Case 2
Figure 11: For each row, the left side plot is for histogram for Residual lifetimes, middle one is for trace plot and right side denotes autocorrelation plot produced based on generated samples by semi-parametric method for unit 4-10. Histograms and autocorrelation plots are constructed based on the samples at lag size 10 for each unit.
Figure 12: For each row, the left side plot is for histogram for Residual lifetimes, middle one is for trace plot and right side denotes autocorrelation plot produced based on generated samples by semi-parametric method for unit 11-18. Histograms and autocorrelation plots are constructed based on the samples at lag size 10 for each unit.
Figure 13: For each row, the left side plot is for histogram for Residual lifetimes, middle one is for trace plot and right side denotes autocorrelation plot produced based on generated samples by semi-parametric method for unit 19-26. Histograms and autocorrelation plots are constructed based on the samples at lag size 10 for each unit.
Figure 14: For each row, the left side plot is for histogram for Residual lifetimes, middle one is for trace plot and right side denotes autocorrelation plot produced based on generated samples by semi-parametric method for unit 27-30. Histograms and autocorrelation plots are constructed based on the samples at lag size 10 for each unit.
A.3 Fatigue - Crack Size Data
Figure 15: For each row, the left side plot is for histogram for Residual lifetimes, middle one is for trace plot and right side denotes autocorrelation plot produced based on generated samples by semi-parametric method for unit 4-5. Histograms and autocorrelation plots are constructed based on the samples at lag size 10 for each unit.
Figure 16: For each row, the left side plot is for histogram for Residual lifetimes, middle one is for trace plot and right side denotes autocorrelation plot produced based on generated samples by semi-parametric method for unit 6-13. Histograms and autocorrelation plots are constructed based on the samples at lag size 10 for each unit.
Figure 17: For each row, the left side plot is for histogram for Residual lifetimes, middle one is for trace plot and right side denotes autocorrelation plot produced based on generated samples by semi-parametric method for unit 14-21. Histograms and autocorrelation plots are constructed based on the samples at lag size 10 for each unit.
Appendix B Histogram of Residual lifetime, Monitoring MCMC convergence for Bayesian parametric method
B.1 Case 1
Figure 18: For each row, the left side plot is for histogram for Residual lifetimes, middle one is for trace plot and right side denotes autocorrelation plot produced based on generated samples by parametric method for unit 4-10. Histograms and autocorrelation plots are constructed based on the samples at lag size 10 for each unit.
B.2 Case 2
Figure 19: For each row, the left side plot is for histogram for Residual lifetimes, middle one is for trace plot and right side denotes autocorrelation plot produced based on generated samples by parametric method for unit 4-10. Histograms and autocorrelation plots are constructed based on the samples at lag size 10 for each unit.
Figure 20: For each row, the left side plot is for histogram for Residual lifetimes, middle one is for trace plot and right side denotes autocorrelation plot produced based on generated samples by parametric method for unit 11-18. Histograms and autocorrelation plots are constructed based on the samples at lag size 10 for each unit.
Figure 21: For each row, the left side plot is for histogram for Residual lifetimes, middle one is for trace plot and right side denotes autocorrelation plot produced based on generated samples by parametric method for unit 19-26. Histograms and autocorrelation plots are constructed based on the samples at lag size 10 for each unit.
Figure 22: For each row, the left side plot is for histogram for Residual lifetimes, middle one is for trace plot and right side denotes autocorrelation plot produced based on generated samples by parametric method for unit 27-30. Histograms and autocorrelation plots are constructed based on the samples at lag size 10 for each unit.
B.3 Fatigue - Crack Size Data
Figure 23: For each row, the left side plot is for histogram for Residual lifetimes, middle one is for trace plot and right side denotes autocorrelation plot produced based on generated samples by parametric method for unit 4-5. Histograms and autocorrelation plots are constructed based on the samples at lag size 10 for each unit.
Figure 24: For each row, the left side plot is for histogram for Residual lifetimes, middle one is for trace plot and right side denotes autocorrelation plot produced based on generated samples by parametric method for unit 6-13. Histograms and autocorrelation plots are constructed based on the samples at lag size 10 for each unit.
Figure 25: For each row, the left side plot is for histogram for Residual lifetimes, middle one is for trace plot and right side denotes autocorrelation plot produced based on generated samples by parametric method for unit 14-21. Histograms and autocorrelation plots are constructed based on the samples at lag size 10 for each unit.
Appendix C Monitoring MCMC convergence of model parameters
C.1 Case 1
Figure 26: For each row, first two plots from the left side is trace plot and autocorrelation plot produced by semi-parametric method and last two plots denotes trace plot and autocorrelation plot produced by parametric method based on the generated samples for $\beta_{i}$, $i=1,..,8$ respectively. Autocorrelation plots are constructed based on the samples at lag size 50.
Figure 27: For each row, first two plots from the left side is trace plot and autocorrelation plot produced by semi-parametric method and last two plots denotes trace plot and autocorrelation plot produced by parametric method based on the generated samples for $\beta_{9},\beta_{10},\alpha,\sigma_{\epsilon}^{2}$ respectively. Autocorrelation plots are constructed based on the samples at lag size 50.
C.1.1 Case 2
Figure 28: For each row, first two plots from the left side is trace plot and autocorrelation plot produced by semi-parametric method and last two plots denotes trace plot and autocorrelation plot produced by parametric method based on the generated samples for $\beta_{i}$, $i=1,..,3$ respectively. Autocorrelation plots are constructed based on the samples at lag size 50.
Figure 29: For each row, first two plots from the left side is trace plot and autocorrelation plot produced by semi-parametric method and last two plots denotes trace plot and autocorrelation plot produced by parametric method based on the generated samples for $\beta_{i}$, $i=4,..,12$ respectively. Autocorrelation plots are constructed based on the samples at lag size 50.
Figure 30: For each row, first two plots from the left side is trace plot and autocorrelation plot produced by semi-parametric method and last two plots denotes trace plot and autocorrelation plot produced by parametric method based on the generated samples for $\beta_{i}$, $i=13,..,21$ respectively. Autocorrelation plots are constructed based on the samples at lag size 50.
Figure 31: For each row, first two plots from the left side is trace plot and autocorrelation plot produced by semi-parametric method and last two plots denotes trace plot and autocorrelation plot produced by parametric method based on the generated samples for $\beta_{i}$, $i=22,..,29$ respectively. Autocorrelation plots are constructed based on the samples at lag size 50.
Figure 32: For each row, first two plots from the left side is trace plot and autocorrelation plot produced by semi-parametric method and last two plots denotes trace plot and autocorrelation plot produced by parametric method based on the generated samples for $\beta_{30},\alpha,\sigma_{\epsilon}^{2}$ respectively. Autocorrelation plots are constructed based on the samples at lag size 50.
C.1.2 Fatigue-Crack Size Data
Figure 33: For each row, first two plots from the left side is trace plot and autocorrelation plot produced by semi-parametric method and last two plots denotes trace plot and autocorrelation plot produced by parametric method based on the generated samples for $\beta_{i}$, $i=1...,4$ respectively. Autocorrelation plots are constructed based on the samples at lag size 50.
Figure 34: For each row, first two plots from the left side is trace plot and autocorrelation plot produced by semi-parametric method and last two plots denotes trace plot and autocorrelation plot produced by parametric method based on the generated samples for $\beta_{i}$, $i=5,..,13$ respectively. Autocorrelation plots are constructed based on the samples at lag size 50.
Figure 35: For each row, first two plots from the left side is trace plot and autocorrelation plot produced by semi-parametric method and last two plots denotes trace plot and autocorrelation plot produced by parametric method based on the generated samples for $\beta_{i}$, $i=14,..,21$ respectively. Autocorrelation plots are constructed based on the samples at lag size 50.
Figure 36: For each row, first two plots from the left side is trace plot and autocorrelation plot produced by semi-parametric method and last two plots denotes trace plot and autocorrelation plot produced by parametric method based on the generated samples for $\alpha,\sigma_{\epsilon}^{2}$ respectively. Autocorrelation plots are constructed based on the samples at lag size 50.
Appendix D Comparison of prediction of error for Parametric and Semi-parametric method
D.1 Case 1
Figure 37: Error produced for predicting residual lifetime. The left and right plots denote prediction error for Bayesian semi-parametric and parametric method for each unit.
D.2 Case 2
Figure 38: Error produced for predicting residual lifetime. The left and right plots denote prediction error for Bayesian semi-parametric and parametric method for each unit.
D.3 Fatigue-Crack Size Data
Figure 39: Error produced for predicting residual lifetime. The left and right plots denote prediction error for Bayesian semi-parametric and parametric method for each unit.